### v2

@@ 3.2 Intuition with two parameters
However, cases at the extreme boundaries can also occur. For instance, if the variance regret $\alpha/\alpha^{\star}$ or the bias regret $\beta/\beta^{\star}$ uniformly dominates the other for one of the experiments, the optimal design minimizes the dominating term. These are desiderable properties as the optimal design always prioritizes the dominating term. Appendix Table 5 provides a complete overview of all possible scenarios. More complex scenarios where researchers may allocate sample sizes to each experiment are possible, see Section 6.
Figure 3: Example for $j=2$ (regret). Different colors correspond to the regret for choosing either experiment. Figure reports $\max\{\alpha(j,\gamma_{j}^{\star})/\alpha^{\star},\beta(j,\gamma_{j}^{\star})/\beta^{\star}\}$ for $j\in\{1,2\}$ as the x–axis parameter varies (columns: $\omega_{2}$, $v_{2}$, $\sigma_{2}$). The vertical dashed line marks indifference, where the two curves intersect; to its left/right, the optimal experiment is the one with lower max regret. Columns vary, respectively, (a) $\omega_{2}$ with $v_{1}=0.5$, $v_{2}=1$, $\sigma_{1}=\sigma_{2}=1$, $\omega_{1}=1$; (b) $v_{2}$ with $\omega=(0.9,1)$, $v_{1}=1$, $\sigma=(1,1)$; (c) $\sigma_{2}$ with $\omega=(0.9,1)$, $v=(1,1)$, $\sigma_{1}=1$.
Example 4 (Numerical illustration).
For an illustrative example, consider choosing $\gamma_{2}^{\star}$ when experiment $j=2$ is twice as precise as $j=1$, with $\sigma_{1}^{2}=\sigma_{2}^{2}=1$, $v_{1}=1$, $v_{2}=0.5$, and $\omega_{1}=1$. The optimal $\gamma_{2}^{\star}$ depends not only on features of design $j$ but also on those of the alternative design j-j, $(v_{-j}^{2},\sigma_{-j}^{2},\omega_{-j})$, through the benchmarks $(\alpha^{\star},\beta^{\star})$. Figure 2 (top row) plots $\alpha(2,\gamma_{2}^{\star})/\alpha^{\star}$ and $\beta(2,\gamma_{2}^{\star})/\beta^{\star}$ at the optimal weight while varying, column by column, $\omega_{2}$, $v_{2}$, and $\sigma_{2}$. The bottom row reports the corresponding $\gamma_{2}^{\star}$. Varying $\omega_{2}$ yields two key effects: both very small and very large $\omega_{2}$ push $\gamma_{2}^{\star}$ toward one, but for different reasons—when $\omega_{2}$ is small the oracle bias decreases faster than the experimental variance (since $\beta^{\star}=\omega_{2}$), while when $\omega_{2}$ is large the oracle bias is $\beta^{\star}=\omega_{1}$ and $\omega_{2}$ has only on the feasible experiment’s bias. Consequently, both regret and the optimal estimator depend on the oracle’s choice of the experiment. In the interior region, the optimal weight first shifts toward the variance–optimal solution (as the oracle bias declines) and then rises again toward a bias–preferred solution. When we vary the experimental and observational variances, we observe the expected patterns: $\gamma_{2}^{\star}$ moves inversely with experimental variance and in the same direction as observational variance. These patterns translate into implications for the regret of the chosen experiment (Figure 3). For the second arm $j=2$, regret decreases sharply as $\omega_{2}$ increases and then decreases, with further decreases becoming gradual once the bias component is already small. By contrast, for the first arm, regret is at best slightly decreasing for small changes in $\omega_{2}$ (when variance effects dominate), but eventually increases with $\omega_{2}$ because the oracle increasingly favors arm $2$, so the opportunity cost of selecting arm $1$ increases. A summary appears in Table 4, with a complete discussion in Appendix C.2. ∎
4 Generalization to moments selection
@@
In many scenarios, we may expect that experiments only identify relevant moments in the data. We now extend the framework to generalized method of moments (GMM) estimators.
@@
Consider a vector of moment conditions $g(\theta)\in\mathbb{R}^{p}$ that stacks all moments (observational and experimental) potentially available for estimation. Let $\Lambda$ denote the Jacobian of the moment function with respect to $\theta$, evaluated at $\theta$ (i.e., $\Lambda:=\partial g(\theta)/\partial\theta^{\top}$). For simplicity, we will assume that $\Lambda$ is known (i.e., moments are linear in $\theta$). This simplifies exposition but it is not necessary in the local asymptotic framework we present in Section 5.1, where $\Lambda$ can be replaced by the Jacobian of $g(\theta)$ evaluated at the observational estimates. For a given covariance matrix $\Sigma$ of the sample moment vector, let $\bar{g}_{\Sigma}$ denote the sample analogue with
@@
where the vector $b$ captures the biases in the moments. The researcher may decide to access only some of the entries of $\bar{g}$ by selecting a subset of experiments due to budget constraints.
Misspecification set.
Let $\mathcal{I}\subseteq\{1,\dots,p\}$ index all potential experimental moments researchers may select (unbiased by design) and $\mathcal{I}^{c}$ the observational ones (potentially misspecified). We posit bounded misspecification
$\mathcal{B}(B)\;=\;\Big\{b\in\mathbb{R}^{p}:\ b_{j}=0\ \text{for }j\in\mathcal{I},\ \ |b_{j}|\leq B\ \text{for }j\in\mathcal{I}^{c}\Big\}.$ (4)
This restriction states that experimental moments have no bias, whereas observational moments have bias bounded by $B$, potentially unknown to the researcher.
@@
where $W$ and $\Sigma$ are chosen by the researcher from a feasible set. The structure of the estimator follows a standard GMM asymptotic expansion (e.g. Andrews et al., 2020). [^7] The weighting matrix $W$ may be singular. The expansion $\hat{\theta}-\theta\;\dot{=}\;-(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,\bar{g}_{\Sigma}$ holds whenever $\Lambda^{\top}W\Lambda$ is nonsingular, which can be forced through the set of constraints.
Here, $W$ may have columns of (effectively) zeros, which excludes the corresponding moments from the estimator. This occurs when researchers faces a trade-off in which experiment to run, leading to different moments $\bar{g}$ observed in the experiment.
The matrix $\Sigma$ depends on the experimental design (e.g., sample–size allocations across experimental moments). The constraints on $W$ and $\Sigma$ effectively limits the class of experiments, sample size allocations and potentially estimators the researcher may consider.
@@
Assumption 3 requires that the variance of the selected (reweighted) moments is uniformly bounded and strictly positive definite (even when $W$ does not select some moments). This ensures that any subset of moments the researcher chooses to use is nondegenerate.
@@
Let $g(\theta)=(g^{\text{obs}}(\theta),\,g^{\text{exp}}_{2}(\theta),\,g^{\text{exp}}_{3}(\theta))^{\top}\in\mathbb{R}^{3}$, where the first component is observational and the other two are experimental. Suppose at most one experimental moment can be used. Selection is encoded by the GMM weight $W\in\mathbb{R}^{3\times 3}$. A simple example of class weighting matrices is
$W^{(2)}=\begin{pmatrix}w_{\mathrm{obs}}&\rho&0\\ \rho&w_{2}&0\\ 0&0&0\end{pmatrix}\qquad W^{(3)}=\begin{pmatrix}w_{\mathrm{obs}}&0&\rho\\ 0&0&0\\ \rho&0&w_{3}\end{pmatrix}$,
provided that $\Lambda^{\top}W\Lambda$ is invertible. ∎
@@
Let the estimand be a linear functional $\tau(\theta)=\omega^{\top}\theta$ with $\omega\in\mathbb{R}^{d}$. Then
$\mathrm{MSE}_{b}(W,\Sigma)$ $=\big(\omega^{\top}\Gamma_{\Lambda}(W)\,b\big)^{2}\;+\;\omega^{\top}\Gamma_{\Lambda}(W)\,\Sigma\,\Gamma_{\Lambda}(W)^{\top}\omega.$ (6)
@@
With a slight abuse of notation, let
$\alpha(W,\Sigma)\;\equiv\;\omega^{\top}\Gamma_{\Lambda}(W)\,\Sigma\,\Gamma_{\Lambda}(W)^{\top}\omega,\qquad\beta(W)\;\equiv\;\big\|\big(\Gamma_{\Lambda}(W)^{\top}\omega\big)_{\mathcal{I}^{c}}\big\|_{1}^{\,2}.$
and $\alpha^{\star},\beta^{\star}$ their corresponding minimizers over $(W,\Sigma)\in\mathcal{D}^{\prime}$. The moment-based variance and bias regret are defined as $\alpha/\alpha^{\star},\beta/\beta^{\star}$ respectively.
The bias now depends on the interaction between the sensitivity $\omega$, the moment Jacobian $\Lambda$ and the identity of which moments are observational and therefore potentially biased. The weighting matrix $W$ determines how these moments influence the downstream estimator $\hat{\theta}$.
As before, (and with a slight abuse of notation) we consider the adaptation regret of the following form:
$\mathcal{R}(W,\Sigma)\;\equiv\;\sup_{B\geq 0}\ \frac{\sup_{b\in\mathcal{B}(B)}\mathrm{MSE}_{b}(W,\Sigma)}{\inf_{(W^{\prime},\Sigma^{\prime})\in\mathcal{D}^{\prime}}\ \mathrm{MSE}_{b}(W^{\prime},\Sigma^{\prime})}.$ (7)
@@
$\mathcal{R}(W,\Sigma)\;=\;\max\!\left\{\,\frac{\alpha(W,\Sigma)}{\alpha^{\star}}\,,\ \frac{\beta(W)}{\beta^{\star}}\,\right\}.$ (8)

### v3

@@ 3.1 Optimal design
However, cases at the extreme boundaries can also occur. If the variance regret $\alpha/\alpha^{\star}$ or the bias regret $\beta/\beta^{\star}$ uniformly dominates the other for one of the experiments, the optimal design minimizes the dominating term. These are desiderable properties as the optimal design always prioritizes the dominating term. Appendix Table 5 provides a complete overview of all possible scenarios. Appendix B.2 provides a numerical example and Section 6 presents a more complex example that include optimally allocating sample size. ∎
3.2 Optimization
Next, we provide an explicit optimization routine that can be solved using off-the-shelf software. We consider observational estimates independent of the experimental ones (but possibly dependent with each other), and experimental estimates independent of each other. This occurs when conducting experiments on independent samples different from the observational sample as in our applications. More general dependences are possible but omitted.
Algorithm 1 Regret-optimal experimental design
1:Inputs: target functional weights $\omega$, variance parameters $(v_{j}^{2})$, observational covariance $\Sigma_{\mathrm{obs}}$, per-unit costs $(c_{j})$, total budget $n$, and feasible experiment set $\mathcal{X}$.
2:Define variance, bias, and sample sizes. For any $(x,\gamma)$ with $x\in\mathcal{X}$ and $0\leq\gamma_{j}\leq 1$, set $s_{j}=x_{j}\gamma_{j}$ and compute:
  • the sample allocation (8), and the definition of $\alpha(s)$ in (9);
  • the bias contribution $\beta(s)$ as defined below (9).
3:Compute oracle quantities. Obtain $\alpha^{\star}$ and $\beta^{\star}$ by solving the two optimization problems, i.e.,
  • minimize $\alpha(s)$ over $(x,\gamma,s)$ subject to $x\in\mathcal{X}$, $x_{j}\in\{0,1\}$, $s_{j}=x_{j}\gamma_{j}$;
  • minimize $\sum_{j}x_{j}|\omega_{j}|$ over $x\in\mathcal{X}$ with $x_{j}\in\{0,1\}$.
4:Solve the minimax MIQP. Solve (10)–(14) with a off-the-shelf solver.
5:Outputs. Let $(x^{\star},\gamma^{\star},s^{\star},t^{\star})$ denote an optimizer of (10)–(14). Report:
  • the regret-optimal experiment set $\mathcal{E}^{\star}=\{j:x_{j}^{\star}=1\}$ and shrinkage weights $\gamma^{\star}$;
  • the implied $s^{\star}$ and regret factor $t^{\star}=\max\{\alpha/\alpha^{\star},\beta/\beta^{\star}\}$;
  • the optimal sample sizes $n_{j}^{\star}$ obtained by plugging $s^{\star}$ into (8).
Notation and decision variables
For given $(\mathcal{E},\Sigma,\gamma)$, denote by $\Sigma_{\mathrm{obs}}$ the submatrix of $\Sigma$ corresponding to the variance–covariance matrix of the observational estimates (possibly non-diagonal), and let $\mathbb{V}(\tilde{\theta}_{j}^{\mathrm{exp}})=v_{j}^{2}/n_{j}$ be the variance of the experimental estimator for component $j$, where $n_{j}$ denotes the sample size allocated to that experiment. Let $x_{j}=1\{j\in\mathcal{E}\}\in\{0,1\}$ indicate whether the experiment identifies component $j$, and let $x\in\mathcal{X}$, where $\mathcal{X}$ denotes the constraint set of feasible experiments. We optimize over $(x,\gamma,n_{j})$ under a total budget constraint $\sum_{j\in\mathcal{E}}c_{j}n_{j}=n,$ where $c_{j}>0$ denotes a per-unit cost for experiment.
Variance, bias and optimal sample size
Under independence of experimental and observational estimators, the variance of $\tau(\theta)$ can be written as
$\sum_{j}\omega_{j}^{2}s_{j}^{2}\frac{v_{j}^{2}}{n_{j}}\;+\;\big(\omega\odot(1-s)\big)^{\top}\Sigma_{\mathrm{obs}}\big(\omega\odot(1-s)\big),\qquad s_{j}\equiv x_{j}\gamma_{j},$ (7)
where $\odot$ denotes the elementwise product. The sample sizes $n_{j}$ enter only through the first (experimental) component. By the standard Neyman allocation with variable costs (Gerber and Green, 2012), the minimizer over $(n_{j})$ for fixed $s$ is given by
$n_{j}^{\star}\;=\;n\,\frac{|\omega_{j}|v_{j}s_{j}/\sqrt{c_{j}}}{\sum_{k}|\omega_{k}|v_{k}s_{k}\sqrt{c_{k}}}.$ (8)
Substituting (8) into the experimental variance and with a slight abuse of notation, let
$\alpha(s)\;\equiv\;\frac{1}{n}\Big(\sum_{j}s_{j}|\omega_{j}|v_{j}\sqrt{c_{j}}\Big)^{2}+\big(\omega\odot(1-s)\big)^{\top}\Sigma_{\mathrm{obs}}\big(\omega\odot(1-s)\big),\quad\beta(s)\;\equiv\;\Big(\sum_{j}(1-s_{j})|\omega_{j}|\Big)^{2},$ (9)
the variance and bias contributions.
Oracle solutions
The next step is to compute $\alpha^{\star}$ and $\beta^{\star}$. These solve
$\alpha^{\star}=\min_{x,\gamma,s}\alpha(s),$ $\text{s.t. }0\leq\gamma_{j}\leq 1,\;\;x_{j}\in\{0,1\},\;\;x\in\mathcal{X},$
$0\leq s_{j}\leq\gamma_{j},\;\;s_{j}\leq x_{j},\;\;s_{j}\geq\gamma_{j}+x_{j}-1,\quad\forall j,$
$\sqrt{\beta^{\star}}=\min_{x^{\prime}}\sum_{j}(1-x_{j}^{\prime})|\omega_{j}|,$ $\text{s.t. }x^{\prime}\in\mathcal{X},\;\;x_{j}^{\prime}\in\{0,1\},\quad\forall j.$
The constraints for $s$ are standard inequalities enforcing $s_{j}=\gamma_{j}x_{j}$ when $x_{j}\in\{0,1\}$ and $0\leq\gamma_{j}\leq 1$. Because $\alpha(s)$ is convex quadratic and $\mathcal{X}$ typically admits a linear (or quadratic) description, the first problem is a mixed-integer convex quadratic program, while the second is a mixed-integer linear program. Both can be solved using off-the-shelf solvers.
Regret-optimal solution
We are now ready to solve for the regret-optimal design. By Theorem 1, we want to minimize $\max\{\alpha/\alpha^{\star},\beta/\beta^{\star}\}$. The corresponding optimization program can be written as
$\min_{x,\gamma,s,t}\quad$ $t$ (10)
s.t. $\alpha(s)\;\leq\;t\,\alpha^{\star},$ (11)
$\beta(s)\;\leq\;t\,\beta^{\star},$ (12)
$0\leq\gamma_{j}\leq 1,\quad x_{j}\in\{0,1\},\quad j=1,\dots,p,$ (13)
$0\leq s_{j}\leq\gamma_{j},\quad s_{j}\leq x_{j},\quad s_{j}\geq\gamma_{j}+x_{j}-1,\quad j=1,\dots,p,$ (14)
$x\in\mathcal{X}\quad\text{(constraint on feasible experiments).}$ (15)
The constraints (14) again enforce $s_{j}=x_{j}\gamma_{j}$ for $x_{j}\in\{0,1\}$ and $0\leq\gamma_{j}\leq 1$. The variable $t$ satisfies $t\geq\alpha/\alpha^{\star}$ and $t\geq\beta/\beta^{\star}$, so at the optimum we have $t=\max\{\alpha/\alpha^{\star},\beta/\beta^{\star}\}$ as desired. Because both $\alpha(s)$ and $\beta(s)$ are convex quadratic functions, the optimization problem (10)–(14) is a mixed-integer quadratic program (MIQP) with linear objective and convex quadratic constraints, and can be passed to standard MIQP solvers.
Remark 4 (“Bias-aware” Neyman allocation).
The optimal sample size solution in Equation (8) has the desirable feature that, conditional on the chosen experiment and shrinkage weights, it takes the form of a Neyman (variance-optimal) allocation. However, because both the experiment and the estimator are selected to trade off variance and worst-case bias, the resulting allocation generally differs from the standard Neyman allocation one would obtain by ignoring the sensitivity of the estimand to misspecification (and the fixed-cost constraints). It is therefore natural to refer to our sample size allocation as a (novel) “bias-aware” Neyman allocation, where the choice of the experiment and shrinkage weights trade-off bias and variance, and optimal sample size minimizes the variance conditional on these choices. ∎
4 General framework for moment selection
@@
In many scenarios, we may expect that experiments only identify relevant moments in the data, and observational estimates are often based on model-implied moment conditions. We now extend the framework to generalized method of moments (GMM) estimators. In addition, we allow for a more general class of ambiguity sets and multivalued estimands.
@@
Consider a vector of moment conditions $g(\theta)\in\mathbb{R}^{p}$ that stacks the “menu of moments” (observational and experimental), i.e. all moments potentially available for estimation. Let $\Lambda$ denote the Jacobian of the moment function with respect to $\theta$, evaluated at $\theta$ (i.e., $\Lambda:=\partial g(\theta)/\partial\theta^{\top}$). For simplicity, we will assume that $\Lambda$ is known (i.e., moments are linear in $\theta$). This simplifies exposition but it is not necessary in the local asymptotic framework we present in Section 5.3, where $\Lambda$ can be replaced by the Jacobian of $g(\theta)$ evaluated at the observational estimates. Let $\bar{g}_{\Sigma}$ denote the sample analogue with
@@
where the vector $b$ captures the biases in the moments. The researcher may access only some of the entries of $\bar{g}$ by selecting a subset of experiments due to budget constraints.
General misspecification set.
Let $\mathcal{I}\subseteq\{1,\dots,p\}$ index all potential experimental moments researchers may select (unbiased by design) and $\mathcal{I}^{c}$ the observational ones (potentially misspecified). We posit bounded misspecification under arbitrary norms
[…]
where the ambiguity set is defined relative to an arbitrary norm $||\cdot||_{l}$. Examples include the $\ell_{1},\ell_{2},\ell_{\infty}$ norms as special cases, as well as norms that may weight observational biases differently, or may even assume no bias for some observational moments (see Example 7).
@@
where $W$ and $\Sigma$ are chosen by the researcher from a feasible set. The structure of the estimator follows a standard GMM asymptotic expansion (e.g. Andrews et al., 2020, 2017). [^11] The weighting matrix $W$ may be singular. The expansion $\hat{\theta}-\theta\;\dot{=}\;-(\Lambda^{\top}W\Lambda)^{-1}\Lambda^{\top}W\,\bar{g}_{\Sigma}$ holds whenever $\Lambda^{\top}W\Lambda$ is nonsingular, which can be forced through the set of constraints in Assumption 3.
Here, $W$ may have columns of zeros, which excludes the corresponding moments from the estimator. This occurs when researchers face a trade-off in which experiment to run, leading to different moments $\bar{g}$ observed in the experiment.
The matrix $\Sigma$ depends on the experimental design (e.g., sample-size allocations across experimental moments). The constraints on $W$ and $\Sigma$ effectively limit the class of experiments, sample size allocations and estimators the researcher may consider.
@@
Assumption 3 requires that the variance of the selected (reweighted) moments is uniformly bounded and strictly positive definite (even when $W$ does not select some moments). This ensures that any subset of moments the researcher chooses to use is nondegenerate. Note that for now we are abstracting from settings with incomplete models studied in Section 5.1.
@@
Let $\theta\in\mathbb{R}^{2}$ and
$g(\theta)=\big(g_{1}^{\text{obs}}(\theta),\,g_{2}^{\text{obs}}(\theta),\,g^{\text{exp}}_{1}(\theta),\,g^{\text{exp}}_{2}(\theta)\big)^{\top}\in\mathbb{R}^{4},$
where the first two components are observational and the last two are experimental. Suppose at most one experimental moment can be used. Selection is encoded by the GMM weight matrix $W\in\mathbb{R}^{4\times 4}$. A simple class of weighting matrices is
$W^{(1)}=\begin{pmatrix}w_{\mathrm{obs},1}&\rho&0&0\\ \rho&w_{\mathrm{obs},2}&0&0\\ 0&0&w_{\mathrm{exp},1}&0\\ 0&0&0&0\end{pmatrix}\qquad W^{(2)}=\begin{pmatrix}w_{\mathrm{obs},1}&\rho&0&0\\ \rho&w_{\mathrm{obs},2}&0&0\\ 0&0&0&0\\ 0&0&0&w_{\mathrm{exp},2}\end{pmatrix}$,
where $W^{(1)}$ selects the second experimental moment $g^{\text{exp}}_{1}$ and sets the weight on $g^{\text{exp}}_{2}$ to zero, while $W^{(2)}$ selects $g^{\text{exp}}_{2}$ and excludes $g^{\text{exp}}_{1}$. ∎
@@
Let the estimand be potentially multivalued $\tau(\theta)\in\mathbb{R}^{q}$, where $\tau(\theta)=\Omega\theta$ with $\Omega\in\mathbb{R}^{q\times d}$ with $\Omega$ different from the zero matrix. Our objective is the sum of MSE for each entry of $\tau$, that is equal to
[…]
@@
[…]
$\alpha_{\Omega}(W,\Sigma)\;\equiv\;\mathrm{Trace}\left(\Omega\Gamma_{\Lambda}(W)\,\Sigma\,\Gamma_{\Lambda}(W)^{\top}\Omega^{\top}\right),\qquad\beta_{l,\Omega}(W)\;\equiv\;\big\|[\Omega\Gamma_{\Lambda}(W)]_{.,\mathcal{I}^{c}}\big\|_{l,*}^{\,2}.$
Let $\alpha_{\Omega}^{\star},\beta_{l,\Omega}^{\star}$ their corresponding minimizers over $(W,\Sigma)\in\mathcal{D}^{\prime}$. The moment-based variance and bias regret are defined as $\alpha_{\Omega}/\alpha_{\Omega}^{\star},\beta_{l,\Omega}/\beta_{l,\Omega}^{\star}$ respectively.
The bias now depends on the interaction between the sensitivity $\Omega$, the moment Jacobian $\Lambda$ and the identity of which moments are observational and therefore potentially biased. For unidimensional $\tau(\theta)$ the bias corresponds to the squared dual norm of $\Omega\Gamma_{\Lambda}(W)$ with respect to $||\cdot||_{l}$. The weighting matrix $W$ determines how these moments influence $\hat{\theta}$.
As before, we consider the adaptation regret with respect to an oracle that knows $B$:
$\mathcal{R}_{l}(W,\Sigma)\;\equiv\;\sup_{B\geq 0}\ \frac{\sup_{b\in\mathcal{B}_{l}(B)}\mathrm{MSE}_{b}(W,\Sigma)}{\inf_{(W^{\prime},\Sigma^{\prime})\in\mathcal{D}^{\prime}}\sup_{b^{\prime}\in\mathcal{B}_{l}(B)}\ \mathrm{MSE}_{b^{\prime}}(W^{\prime},\Sigma^{\prime})}.$ (19)
@@
$\mathcal{R}_{l}(W,\Sigma)\;=\;\max\!\left\{\,\frac{\alpha_{\Omega}(W,\Sigma)}{\alpha_{\Omega}^{\star}}\,,\ \frac{\beta_{l,\Omega}(W)}{\beta_{l,\Omega}^{\star}}\,\right\}.$ (20)